OpenStudy (anonymous):

Motion Problem: At 4:00 Gigi and Madel left the music house. Gigi walked east and reached her house at 5:00. Madel rode her bike west and reached her house at 4:15. Madel traveled twice as fast as Gigi. Their houses are 9km apart. How far didi each travel?

OpenStudy (anonymous):

Gigi travelled 6km, Madel travelled 3km.

OpenStudy (anonymous):

thanks, will you help me in representing? pls?

OpenStudy (anonymous):

Distance = time*speed Gigi Time = 60 Madel Time = 15 Gigi Speed = x Madel speed = 2x Gigi Distance = time*speed = 60*x = 60x Madel Distance = time*speed = 15*2x = 30x 60x + 30x = 9 90x = 9 x = 1/10 Gigi Distance = 60*1/10 = 6km Madel Distance = 15*2*1/10 = 3km

OpenStudy (anonymous):

let gigi travel x distance, and madel y distance.. since east and west are opposite.. x+y=9 Next.. gigi takes 1 hours to travel x , and madel 15 mins to travel y Next.. Sgigig = x/1 and Smadel = y/15 and Smadel = 2*Sgigi

OpenStudy (anonymous):

sorry thats 60 not 1 :D

OpenStudy (anonymous):

THANKS TO YOU BOTH ! its a great help !!!!! God Bless !

OpenStudy (anonymous):

last question, does "*" means multiply? im really slow

OpenStudy (anonymous):

Yes, with x being a constant :)